Under the guidance of a good manager, a monopolistically competitive firm spends an optimal amount of its revenues on advertising. Last month the firm spent $2,000 of its $10,000 revenues on advertising. If its advertising elasticity was 0.5, what was this firm’s price elasticity demand? Does this make the firm’s demand elastic, unit elastic or inelastic?
Answer #1

Advertising expenditure / Revenue = - Advertising elasticity / price elasticity

2000 / 10000 = - 0.5 / price elasticity

price elasticity = -0.5 / 0.2 = -2.5

As it is negative, hence demand is inelastic